OpenStudy (exceptionlearner):

A company that produces DVDs uses the formula C=90n1/3+350 to calculate the cost C in dollars of producing n DVDS per day. How many DVDS can be produced for a cost of $800? A. 15 B. 45 C. 75 D. 125

OpenStudy (johan14th):

Is it C=90n(1/3)+350?

OpenStudy (johan14th):

800=90n(1/3)+350 -350 -350 450=90n(1/3) 450=30n ÷30 ÷30 450/30 =n ?=n